OpenStudy (anonymous):

-2x=2y=10 okay so i understand that i have to get it to y=mx+b. just add 2 to both sides. but the equation after that tells me to divide as in (2y=2x+10' all being divided by 2) with the answer coming out as y=x+5. but what i dont get is how when i added 2 to both sides why didnt the equation change to x+2y=12 ?

OpenStudy (texaschic101):

2x = 2y + 10 ....Is this the original problem ?

OpenStudy (anonymous):

-2x +2y=10

OpenStudy (texaschic101):

ok...-2x + 2y = 10 and you have to change it to y = mx + b. This means you have to get 2y on one side.....first you can add 2x to both sides. -2x + 2y + 2x = 10 + 2x 2y = 2x + 10 Then you would have to divide by 2 to get y by itself. 2y/2 = 2x/2 + 10/2 y = x + 5 I see what you did....the 2 is multiplied by x so adding just the 2, will not get x by itself. You have to add 2x to cancel them on the right side and move them to the left side. Now if the problem was -2 + x + 2y = 10, then you could add two, but when your multiplying, then you need to divide (do the opposite) opposite of multiplication is division and opposite of addition is subtraction and opposite of subtraction is addition. Do you understand ?

OpenStudy (anonymous):

yess! thank you so much !

OpenStudy (texaschic101):

glad to help :)